# 3D-GP-LMVIC: Learning-based Multi-View Image Coding with 3D Gaussian Geometric Priors

### v1

@@ 3.2.1 Image Compression Model
As depicted in Figure 3, the compression and decompression of the image $\bm{x}_{n}$ utilize references from the previously reconstructed image $\bm{\hat{x}}_{n-1}$, the intermediate feature set $\{\bm{f}^{i}_{n-1}\mid i=1,2,3\}$ obtained during the reconstruction of $\bm{\hat{x}}_{n-1}$, and the reconstructed depth maps $\bm{\hat{d}}_{n-1}$ and $\bm{\hat{d}}_{n}$. The image $\bm{x}_{n}$ is encoded into $\bm{y}_{n}$ by the image encoder $I_{\text{E}}$. Subsequently, $\bm{y}_{n}$ is quantized into $\bm{\hat{y}}_{n}$, which is then decoded by the image decoder $I_{\text{D}}$ to produce the reconstructed image $\bm{\hat{x}}_{n}$. This process can be formalized as:
$\bm{y}_{n}$ $=I_{\text{E}}(\bm{x}_{n},\bm{\hat{x}}_{n-1},\{\bm{f}^{i}_{n-1}\mid i=1,2,3\},\bm{\hat{d}}_{n-1},\bm{\hat{d}}_{n}),$ (5)
$\bm{\hat{y}}_{n}$ $=Q(\bm{y}_{n}),$
$\bm{\hat{x}}_{n}$ $=I_{\text{D}}(\bm{\hat{y}}_{n},\bm{\hat{x}}_{n-1},\{\bm{f}^{i}_{n-1}\mid i=1,2,3\},\bm{\hat{d}}_{n-1},\bm{\hat{d}}_{n}).$
For entropy coding, we employ the hyperprior entropy model (Ballé et al., 2018) and the quadtree partition-based entropy model (QPEM) (Li et al., 2023). The hyperprior entropy model transforms $\bm{y}_{n}$ into a hyperprior representation $\bm{z}_{n}$, which facilitates the accurate modeling of the probability distribution of $\bm{\hat{y}}_{n}$. QPEM introduces diverse spatial contexts for enhanced entropy coding efficiency while ensuring fast coding speed.
@@
Disparity extractor. As illustrated by the brown dotted-dashed line box in Figure 3, we firstly employ the disparity estimation (DPE) module to derive the disparity map $\bm{\Delta}_{n}$ and the corresponding mask $\bm{x}_{n,\text{m}}$, following the method outlined in Section 3.1, using $\bm{\hat{d}}_{n-1}$ and $\bm{\hat{d}}_{n}$. Subsequently, $\bm{\Delta}_{n}$ undergoes a series of downsampling operations to produce multi-scale disparity maps $\{\bm{\Delta}^{i}_{n}\mid i=1,2,3\}$, which will facilitate multi-scale feature alignment. The mask $\bm{x}_{n,\text{m}}$ is further processed by the disparity mask extractor to extract feature masks $\{\bm{f}^{i}_{n,\text{m}}\mid i=1,2,3,4\}$, which will aid in isolating relevant features.
@@
Reference feature extractor. The reference feature extractor takes as inputs $\bm{\hat{x}}_{n-1}$, $\{\bm{f}^{i}_{n-1}\mid i=1,2,3\}$, and $\bm{\Delta}^{3}_{n}$ to extract multi-scale reference features $\{\bm{h}^{i}_{n-1}\mid i=1,2,3,4\}$, as shown by the light blue dotted-dashed line box in Figure 3. Specifically, $\bm{h}^{4}_{n-1}$ is derived by aligning $\bm{h}^{3}_{n-1}$ with $\bm{\Delta}^{3}_{n}$ and subsequently applying a convolutional layer.
Image context transfer module. To incorporate the reference feature $\{\bm{h}^{i}_{n-1}\mid i=1,2,3\}$ obtained from the $(n-1)$-th view into the image backbone encoder and decoder, enhancing feature representation, we introduce the image context transfer (ICT) module. As depicted by the orange-red dotted-dashed line box in Figure 3, the input feature $\bm{f}^{i^{*}}_{n}$ from the backbone network and the aligned feature of $\bm{h}^{i}_{n-1}$ via $\bm{\Delta}^{i}_{n}$ are concatenated after a convolutional layer. The concatenated feature is then element-wise multiplied with $\bm{f}^{i}_{n,\text{m}}$, followed by a residual block for feature refinement. The final step involves element-wise addition of this refined feature to $\bm{f}^{i^{*}}_{n}$, yielding the output feature $\bm{f}^{i}_{n}$.
@@ 3.2.2 Depth Map Compression Model
The compression and decompression of depth map $\bm{d}_{n}$ utilize $\bm{\hat{d}}_{n-1}$ as a reference. The depth map $\bm{d}_{n}$ is processed by the depth encoder $D_{\text{E}}$ to produce the latent representation $\bm{y}_{d_{n}}$, which is subsequently quantized into $\bm{\hat{y}}_{d_{n}}$. The quantized representation is decoded by the depth decoder $D_{\text{D}}$ to reconstruct the depth map $\bm{\hat{d}}_{n}$. This process is formalized as follows:
$\bm{y}_{d_{n}}$ $=D_{\text{E}}(\bm{d}_{n},\bm{\hat{d}}_{n-1}),$ (6)
$\bm{\hat{y}}_{d_{n}}$ $=Q(\bm{y}_{d_{n}}),$
$\bm{\hat{d}}_{n}$ $=D_{\text{D}}(\bm{\hat{y}}_{d_{n}},\bm{\hat{d}}_{n-1}).$
The entropy coding scheme also incorporates the hyperprior entropy model and the QPEM. The latent representation $\bm{y}_{d_{n}}$ is transformed into a hyperprior representation $\bm{z}_{d_{n}}$ using the hyperprior entropy model. Additional details of the depth map compression model are provided in Appendix A.
@@ 3.2.3 Training Loss
For each training step, a randomly selected subsequence of length 4 from a multi-view sequence serves as the training sample. The training loss comprises the distortion losses for both the reconstructed image and depth map, as well as the estimated compression rates for the encoded image and depth map, for each view in the training sample:
$L=\sum_{n=s}^{s+3}w_{n-s+1}\left[\lambda_{\text{img}}D(\bm{x}_{n},\bm{\hat{x}}_{n})+\lambda_{\text{dep}}\text{MSE}(\bm{d}_{n},\bm{\hat{d}}_{n})+R(\bm{\hat{y}}_{n})+R(\bm{\hat{z}}_{n})+R(\bm{\hat{y}}_{d_{n}})+R(\bm{\hat{z}}_{d_{n}})\right],$ (7)
where $D(\cdot,\cdot)$ denotes the distortion, $\text{MSE}(\cdot,\cdot)$ represents the mean squared error (MSE), and $R(\cdot)$ indicates the estimated compression rates. The hyperparameters $\lambda_{\text{img}}$ and $\lambda_{\text{dep}}$ control the contributions of the image and depth map distortion losses, respectively. The weights $\{w_{i}\mid i=1,2,3,4\}$, as referenced from Li et al. (2023), adjust the influence of each view on the overall training loss.
3.3 Multi-View Sequence Ordering
Given the significant impact of field-of-view overlap between adjacent views on inter-view correlations, we propose a multi-view sequence ordering method to alleviate the issue of insufficient overlap in certain sequences. We define a distance metric to evaluate inter-view overlap and employ a greedy algorithm to find an improved sequence.
In Eq. 3, if $V_{n-1}V_{n}^{-1}=I$, then $(x_{n},y_{n})=(x_{n-1},y_{n-1})$. This indicates that each pixel in the $n$-th view lies within the valid image area of the $(n-1)$-th view, indicating high overlap. Thus, for any two views $i$ and $j$, we measure overlap by the proximity of $V_{i}V_{j}^{-1}$ to the identity matrix:
$\mathcal{D}_{\mathcal{V}}(i,j)=\|V_{i}V_{j}^{-1}-I\|.$ (8)
Appendix B demonstrates that $\mathcal{D}_{\mathcal{V}}(i,j)$ is a distance metric for both the 2-norm and Frobenius norm. The Frobenius norm is utilized in our experiments. After determining pairwise distances, a greedy algorithm is employed, starting from an initial sequence with only one view, iteratively selecting the view closest to the last view in the sequence.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our model on three multi-view image datasets: Tanks&Temples (Knapitsch et al., 2017), Mip-NeRF 360 (Barron et al., 2022), and Deep Blending (Hedman et al., 2018). These datasets feature a wide variety of indoor and outdoor scenes, each containing dozens to over a thousand images captured from different views. Further details on the datasets are provided in Appendix C.
@@ 4.1 Experimental Setup
3D-GP-LMVIC -47.48% -63.69% -34.69% -40.25% -27.31% -54.15%
Benchmarks. We compare our approach against several baselines, including traditional multi-view codec: MV-HEVC (Hannuksela et al., 2015); learning-based multi-view image codecs: two variants of HESIC (Deng et al., 2021), MASIC (Deng et al., 2023), SASIC (Wödlinger et al., 2022), two variants of LDMIC (Zhang et al., 2023), and two variants of BiSIC (Liu et al., 2024); as well as the 3D-GS compression method: HAC (Chen et al., 2024). Further details on the baseline configurations are provided in Appendix C.
@@
Implementation Details. The model was trained using five different configurations of ($\lambda_{\text{img}}$, $\lambda_{\text{dep}}$): $\left((256,64),(512,128),(1024,128),(2048,128),(4096,128)\right)$ when the image distortion loss is MSE, and $\left((8,64),(16,128),(32,128),(64,128),(128,128)\right)$ when using MS-SSIM. The weights $w_{i}$ for four consecutive views were set to $(0.5,1.2,0.5,0.9)$. The model was trained for 300 epochs with an initial learning rate of $10^{-4}$, which was progressively decayed by a factor of 0.5 every 60 epochs.
@@ 4.2 Experimental Results
Coding performance. Figure 4 presents the rate-distortion curves of the compared methods, while Table 1 summarizes the BDBR of each codec relative to MV-HEVC. Across the three datasets, the proposed 3D-GP-LMVIC consistently outperforms the baselines in both PSNR and MS-SSIM, demonstrating its effectiveness in reducing inter-view redundancy. For instance, on the Tanks&Temples dataset, 3D-GP-LMVIC achieves a BDBR reduction of 16.59% for PSNR and 13.73% for MS-SSIM compared to BiSIC. The BDBR of HAC is relatively higher, likely due to the inclusion of 3D scene information in addition to 2D image representations.
Encoding and decoding time comparison. Table 2 presents the encoding and decoding runtimes of six learning-based image codecs, evaluated on a platform equipped with an Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz and a GPU with 10,752 parallel processing cores. The neural network components of the codecs were executed on the GPU, while the entropy coding was handled by the CPU. 3D-GP-LMVIC achieved encoding and decoding times of 0.19s and 0.18s, respectively, making it one of the faster methods. This speed can be attributed to its relatively simple network structure and the QPEM, which balances coding speed and compression efficiency. Both SASIC and LDMIC-Fast showed fast runtimes due to their use of highly parallelizable hyperprior and checkerboard entropy models, respectively. Besides, HESIC+, MASIC, and LDMIC utilized autoregressive entropy models, resulting in comparatively slower runtimes.
Table 2: Average encoding and decoding runtime of learning-based codecs on Tanks&Temples (image resolution: $978\times 546$).
Operation Codecs
HESIC+ MASIC SASIC LDMIC-Fast LDMIC 3D-GP-LMVIC
Encoding 4.35s 4.38s 0.06s 0.11s 4.24s 0.19s
Decoding 10.73s 10.78s 0.09s 0.09s 10.63s 0.18s
@@ 4.3 Ablation Study
Codec components. To assess the contribution of codec components, we performed ablation experiments on the Tanks&Temples dataset. The rate-distortion curves are shown in Figure 5. Specifically, we evaluated the following baselines: (1) Separate: encoding and decoding without cross-view information; (2) Concatenation: direct feature concatenation from reference view without alignment; (3) W/O Mask: removal of both image and depth mask; (4) W/O Dep.Pred: excluding depth prediction in the depth map compression model. These baselines resulted in bitrate increases of 41.07% (44.24%), 42.75% (47.52%), 7.19% (8.47%), and 8.03% (8.02%) for PSNR (MS-SSIM), respectively, compared to the proposed method. The experimental results validate the effectiveness of the proposed components. Appendix D provides further analysis on alignment.

### v2

@@ 3.2.1 Image Compression Model
As shown in Figure 3, the disparity extractor $DISE$ utilizes reconstructed depth maps $\boldsymbol{\hat{d}}_{n-1}$ and $\boldsymbol{\hat{d}}_{n}$ to extract multi-scale disparities and feature masks. The reference feature extractor $RFE$ generates multi-scale reference features from the reconstructed image $\boldsymbol{\hat{x}}_{n-1}$ and its intermediate reconstruction features $\{\boldsymbol{f}^{i}_{n-1}\mid i=1,2,3\}$. Subsequently, the image encoder $IE$ and decoder $ID$ incorporate the reference features, aligned using the extracted disparities, into the backbone network. This process is formalized as:
$\boldsymbol{y}_{n}$ $=IE(\boldsymbol{x}_{n},DISE(\boldsymbol{\hat{d}}_{n-1},\boldsymbol{\hat{d}}_{n}),RFE(\boldsymbol{\hat{x}}_{n-1},\{\boldsymbol{f}^{i}_{n-1}\})),$ (5)
$\boldsymbol{\hat{y}}_{n}$ $=Q(\boldsymbol{y}_{n}),$
$\boldsymbol{\hat{x}}_{n}$ $=ID(\boldsymbol{\hat{y}}_{n},DISE(\boldsymbol{\hat{d}}_{n-1},\boldsymbol{\hat{d}}_{n}),RFE(\boldsymbol{\hat{x}}_{n-1},\{\boldsymbol{f}^{i}_{n-1}\})).$
For entropy coding, we utilize the hyperprior entropy model (Ballé et al., 2018) and the quadtree partition-based entropy model (QPEM) (Li et al., 2023). The hyperprior entropy model transforms $\boldsymbol{y}_{n}$ into a hyperprior representation $\boldsymbol{z}_{n}$. The quantized hyperprior representation $\boldsymbol{\hat{z}}_{n}$ is then used to accurately model the probability distribution of $\boldsymbol{\hat{y}}_{n}$. The conditional probability distribution $p_{\boldsymbol{\hat{y}}_{n}|\boldsymbol{\hat{z}}_{n}}$ is defined as:
$p_{\boldsymbol{\hat{y}}_{n}|\boldsymbol{\hat{z}}_{n}}(\boldsymbol{\hat{y}}_{n}|\boldsymbol{\hat{z}}_{n})\sim\mathcal{N}(\boldsymbol{\mu}_{n},\boldsymbol{\sigma}_{n}^{2}).$ (6)
@@
Disparity extractor. As illustrated in Figure 3, we firstly employ the disparity estimation (DPE) module to derive the disparity map $\boldsymbol{\Delta}_{n}$ and the corresponding mask $\boldsymbol{x}_{n,\text{m}}$, using $\boldsymbol{\hat{d}}_{n-1}$ and $\boldsymbol{\hat{d}}_{n}$, following the method outlined in Section 3.1. Subsequently, $\boldsymbol{\Delta}_{n}$ undergoes a series of downsampling operations to produce multi-scale disparity maps $\{\boldsymbol{\Delta}^{i}_{n}\mid i=1,2,3\}$, which will facilitate multi-scale feature alignment. The mask $\boldsymbol{x}_{n,\text{m}}$ is further processed by the disparity mask extractor to extract feature masks $\{\boldsymbol{f}^{i}_{n,\text{m}}\mid i=1,2,3,4\}$.
@@
Reference feature extractor. The reference feature extractor takes $\boldsymbol{\hat{x}}_{n-1}$, $\{\boldsymbol{f}^{i}_{n-1}\mid i=1,2,3\}$, and $\boldsymbol{\Delta}^{3}_{n}$ as inputs to extract multi-scale reference features $\{\boldsymbol{h}^{i}_{n-1}\mid i=1,2,3,4\}$, as shown in Figure 3.
Figure 4: Illustration of the proposed image context transfer module.
Image context transfer module. To incorporate the reference feature $\{\boldsymbol{h}^{i}_{n-1}\mid i=1,2,3\}$ obtained from the $(n-1)$-th view into the image backbone encoder and decoder, enhancing feature representation, we introduce the image context transfer (ICT) module. As depicted in Figure 4, the module enhances the input feature $\boldsymbol{f}^{i^{*}}_{n}$ from the backbone network by leveraging the aligned reference feature $\boldsymbol{h}^{i}_{n-1}$ via $\boldsymbol{\Delta}^{i}_{n}$. By applying feature masks, the module filters relevant information and refines the features, ultimately producing the output feature $\boldsymbol{f}^{i}_{n}$ through a residual enhancement process.
@@ 3.2.2 Depth Map Compression Model
The compression and decompression of the depth map $\boldsymbol{d}_{n}$ leverage $\boldsymbol{\hat{d}}_{n-1}$ as a reference. Initially, $\boldsymbol{\hat{d}}_{n-1}$ is processed by the depth prediction extractor $DEPE$, which generates multi-scale depth prediction features and corresponding feature masks. Subsequently, the depth encoder $DE$ and decoder $DD$ integrate these extracted features and masks into the backbone network. This process is formalized as:
$\boldsymbol{y}_{d_{n}}$ $=DE(\boldsymbol{d}_{n},DEPE(\boldsymbol{\hat{d}}_{n-1})),$ (7)
$\boldsymbol{\hat{y}}_{d_{n}}$ $=Q(\boldsymbol{y}_{d_{n}}),$
$\boldsymbol{\hat{d}}_{n}$ $=DD(\boldsymbol{\hat{y}}_{d_{n}},DEPE(\boldsymbol{\hat{d}}_{n-1})).$
The entropy coding scheme incorporates both the hyperprior entropy model and the QPEM. The latent representation $\boldsymbol{y}_{d_{n}}$ is transformed into a hyperprior representation $\boldsymbol{z}_{d_{n}}$ using the hyperprior entropy model. Similar to the image compression model, the quantized hyperprior representation $\boldsymbol{\hat{z}}_{d_{n}}$ is used to model the probability distribution of $\boldsymbol{\hat{y}}_{d_{n}}$. Additional details about the depth map compression model are provided in Appendix B.
3.2.3 Multi-View Sequence Ordering
Given the significant impact of FoV overlap between adjacent views on inter-view correlations, we propose a multi-view sequence ordering method to alleviate the issue of insufficient overlap in unordered sequences. We define a distance metric to evaluate inter-view overlap and employ a greedy algorithm to find an improved sequence.
In (3), if $V_{n-1}V_{n}^{-1}=I$, then $(x_{n},y_{n})=(x_{n-1},y_{n-1})$. This indicates that each pixel in the $n$-th view lies within the valid image area of the $(n-1)$-th view, indicating high overlap. Thus, for any two views $i$ and $j$, we measure overlap by the proximity of $V_{i}V_{j}^{-1}$ to the identity matrix:
$\mathcal{D}_{\mathcal{V}}(i,j)=\|V_{i}V_{j}^{-1}-I\|.$ (8)
Appendix C proves that $\mathcal{D}_{\mathcal{V}}(i,j)$ is a distance metric for both the 2-norm and Frobenius norm. The Frobenius norm is utilized in our experiments. After determining pairwise distances, a greedy algorithm is employed, starting from an initial sequence with only one view, iteratively selecting the view closest to the last view in the sequence.
@@ 3.2.4 Training Loss
For each training step, a randomly selected subsequence of length 4 from a multi-view sequence serves as the training sample. The training loss comprises the distortion losses for both the reconstructed image and depth map, as well as the estimated compression rates for the encoded image and depth map:
$L=\sum_{n=s}^{s+3}w_{n-s+1}\Big{[}\lambda_{\text{img}}D(\boldsymbol{x}_{n},\boldsymbol{\hat{x}}_{n})+\lambda_{\text{dep}}\text{MSE}(\boldsymbol{d}_{n},\boldsymbol{\hat{d}}_{n})$ (9)
$+R(\boldsymbol{\hat{y}}_{n})+R(\boldsymbol{\hat{z}}_{n})+R(\boldsymbol{\hat{y}}_{d_{n}})+R(\boldsymbol{\hat{z}}_{d_{n}})\Big{]},$
where $D(\cdot,\cdot)$ denotes the distortion, $\text{MSE}(\cdot,\cdot)$ represents the mean squared error (MSE), and $R(\cdot)$ indicates the estimated compression rates. The hyperparameters $\lambda_{\text{img}}$ and $\lambda_{\text{dep}}$ control the contributions of the image and depth map distortion losses, respectively. The weights $\{w_{i}\mid i=1,2,3,4\}$ adjust the influence of each view on the overall training loss.
@@
3D-LMVIC -47.48% -63.69% -34.69% -40.25% -27.31% -54.15%
Table 2: Average alignment quality (PSNR, MS-SSIM) of different alignment methods on the Train scene of the Tanks&Temples dataset.
Metrics Methods
HT PM SPyNet PWC-Net FlowFormer++ 3D-GS COLMAP MVSFormer++ Proposed
PSNR 15.16 17.94 16.12 17.59 18.08 17.36 14.32 15.31 18.14
MS-SSIM 0.5435 0.7633 0.6289 0.7707 0.7863 0.7410 0.7446 0.5544 0.8053
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our model on three multi-view image datasets: Tanks&Temples (Knapitsch et al., 2017), Mip-NeRF 360 (Barron et al., 2022), and Deep Blending (Hedman et al., 2018). These datasets feature a wide variety of indoor and outdoor scenes, each containing dozens to over a thousand images captured from different views. Further details on the datasets are provided in Appendix D.
Benchmarks. We compare our approach against several baselines, including traditional multi-view codec: MV-HEVC (Hannuksela et al., 2015); learning-based multi-view image codecs: two variants of HESIC (Deng et al., 2021), MASIC (Deng et al., 2023), SASIC (Wödlinger et al., 2022), two variants of LDMIC (Zhang et al., 2023), and two variants of BiSIC (Liu et al., 2024); as well as the 3D-GS compression method: HAC (Chen et al., 2024). Further details on the baseline configurations are provided in Appendix D.
@@ 4.1 Experimental Setup
Implementation Details. The model was trained using five different configurations of ($\lambda_{\text{img}}$, $\lambda_{\text{dep}}$): $\left((256,64),(512,128),(1024,128),(2048,128),(4096,128)\right)$ when the image distortion loss is MSE, and $\left((8,64),(16,128),(32,128),(64,128),(128,128)\right)$ when using MS-SSIM. The weights $w_{i}$ for four consecutive views were set to $(0.5,1.2,0.5,0.9)$ as referenced from Li et al. (2023). The model was trained for 300 epochs with an initial learning rate of $10^{-4}$, which was progressively decayed by a factor of 0.5 every 60 epochs.
@@ 4.2 Experimental Results
Coding performance. Figure 5 presents the rate-distortion curves of the compared methods, while Table 1 summarizes the BDBR of each codec relative to MV-HEVC. Across the three datasets, the proposed 3D-LMVIC consistently outperforms the baselines in both PSNR and MS-SSIM, demonstrating its effectiveness in reducing inter-view redundancy. For instance, on the Tanks&Temples dataset, 3D-LMVIC achieves a BDBR reduction of 16.59% for PSNR and 13.73% for MS-SSIM compared to BiSIC. The BDBR of HAC is relatively higher, likely due to the inclusion of 3D scene information in addition to 2D image representations. Appendix F provides examples of visual comparisons. Appendix G presents an analysis of computational complexity. Appendix H includes supplementary experiments on coding performance.
Reference View
Target View
HT
12.82/0.6065
PWC-Net
9.65/0.2840
PM
10.05/0.4078
FlowFormer++
13.94/0.7216
SPyNet
10.25/0.1990
Proposed
13.94/0.7217
4.3 Alignment Experiments
To evaluate the effectiveness of the proposed 3D Gaussian geometric priors-based alignment method, we conducted alignment experiments on the Train scene from the Tanks&Temples dataset. The baselines for comparison include:
1. Alignment methods commonly used in learning-based multi-view image codecs, such as Homography Transformation (HT) (Deng et al., 2021) and Patch Matching (PM) (Huang et al., 2023).
2. Optical flow estimation methods, such as SPyNet (Ranjan & Black, 2017), PWC-Net (Sun et al., 2018), and FlowFormer++ (Shi et al., 2023).
3. Depth map estimation methods, including original 3D-GS (Kerbl et al., 2023), COLMAP (Schonberger & Frahm, 2016; Schönberger et al., 2016) and MVSFormer++ (Chenjie Cao & Fu, 2024).
Alignment quality was assessed by computing PSNR and MS-SSIM between the aligned reference view images and the target view images. Table 2 summarizes the average alignment quality and runtime for each method. The proposed method outperformed the baselines in both PSNR and MS-SSIM, indicating its effectiveness in capturing complex disparities between views. Figure 6 provides visual comparisons, demonstrating that the proposed method achieves closer alignment with the target view images. Appendix D further investigates the relationship between the mask defined in (4) and the ghosting artifacts introduced during alignment.
@@ 4.4 Ablation Study
Codec components. To assess the contribution of codec components, we performed ablation experiments on the Tanks&Temples dataset. The rate-distortion curves are shown in Figure 7. Specifically, we evaluated the following baselines: (1) Separate: encoding and decoding without cross-view information; (2) Concatenation: direct feature concatenation from reference view without alignment; (3) W/O Mask: removal of both image and depth mask; (4) W/O Dep.Pred: excluding depth prediction in the depth map compression model. These baselines resulted in bitrate increases of 41.07% (44.24%), 42.75% (47.52%), 7.19% (8.47%), and 8.03% (8.02%) for PSNR (MS-SSIM), respectively, compared to the proposed method. The experimental results validate the effectiveness of the proposed components.